A plane flies from Entebbe towards Kiko for 596 km on a bearing of 232 degrees. It then flies from Kikoto Daro for 880 km on a bearing of 258 degrees. It finally flies to Haraka for 458 km on a bearing of 346 degrees from Daro.

(a) How far north or south of Entebbe is /15mrks

(i) Kiko (ii) Haraka?

(b) How far west or east of Entebbe is Haraka?

(c) Calculate the bearing of Haraka from Entebbe


1
Expert's answer
2021-09-03T16:11:12-0400
"E(0,0)"

"K(596\sin(232\degree), 596\cos(232\degree))"

"D(596\sin(232\degree)+880\sin(258\degree),"

"596\cos(232\degree)+880\cos(258\degree))"

"H(596\sin(232\degree)+880\sin(258\degree)+458\sin(346\degree),"

"596\cos(232\degree)+880\cos(258\degree)+458\cos(346\degree))"



(a)

(i)

"K(596\sin(232\degree), 596\cos(232\degree))"


"y=596\cos(232\degree)\ km\approx-367\ km"

Kiko is 367 km south of Entebbe.


(ii)

"H(596\sin(232\degree)+880\sin(258\degree)+458\sin(346\degree),"

"596\cos(232\degree)+880\cos(258\degree)+458\cos(346\degree))"


"y=596\cos(232\degree)+880\cos(258\degree)+458\cos(346\degree))\ km"




"\approx-105.5\ km"

Haraka is 105.5 km south of Entebbe.

(b)


"H(596\sin(232\degree)+880\sin(258\degree)+458\sin(346\degree),"

"596\cos(232\degree)+880\cos(258\degree)+458\cos(346\degree))"


"x=596\sin(232\degree)+880\sin(258\degree)+458\sin(346\degree))\ km"




"\approx-1441.2\ km"

Haraka is 1441.2 km west of Entebbe.


(c)


"H(596\sin(232\degree)+880\sin(258\degree)+458\sin(346\degree),"

"596\cos(232\degree)+880\cos(258\degree)+458\cos(346\degree))"


"\tan\theta=\dfrac{y}{x}"

"=\dfrac{596\cos(232\degree)+880\cos(258\degree)+458\cos(346\degree)}{596\sin(232\degree)+880\sin(258\degree)+458\sin(346\degree)}"

"\approx0.0732"

"\theta\approx4.1867\degree"

"bearing=360\degree-90\degree-\theta\approx265.8\degree"

The bearing of Haraka from Entebbe is "265.8\degree ."